OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5.

OpenStudy (tennistar):

first distribute and take the a + 5 out of the parenthesis

OpenStudy (tennistar):

so you get 2a + 10

OpenStudy (apoorvk):

Ohkay, just substitute the given numerical values of 'a' and 'P' on the formula 'P = 2(a+b)', and you will land up with: 15 = 2(a+5) Can you solve this to find 'a'?

OpenStudy (tennistar):

2a + 10 = 15

OpenStudy (tennistar):

now try solving

OpenStudy (tennistar):

subtract 10 from both sides and you get 2a = 5

OpenStudy (tennistar):

then you divide 2 from both sides and you get \[a = 5 / 2\]

OpenStudy (tennistar):

5 over 2

OpenStudy (tennistar):

do you understand now?

OpenStudy (anonymous):

yes